OpenStudy (ibbutibbu.):

<----- Save this poor child from being abused by numbers! Donate knowledge today!

OpenStudy (anonymous):

Do you have a question?

OpenStudy (ibbutibbu.):

PART C ONLYYY

myininaya (myininaya):

So what is the function?

myininaya (myininaya):

f(x)=-x+6

myininaya (myininaya):

where x is the distance from the hoop and y is the numbers of hoops made

myininaya (myininaya):

(x,y) is an ordered pair

myininaya (myininaya):

So if I asked you what is the ordered pair for when he 2 meters away from the hoop, would you be able to find that ordered pair?

myininaya (myininaya):

since y=f(x) you can also view the ordered pair as (x,f(x))

myininaya (myininaya):

So what am I asking really? I'm asking you to give me the ordered pair for when x=2. Could you do that?

OpenStudy (ibbutibbu.):

Ummm is it (2,8)?

myininaya (myininaya):

how did you get 8?

myininaya (myininaya):

Is the function f(x)=-x+6 or f(x)=x+6?

OpenStudy (ibbutibbu.):

the first one

myininaya (myininaya):

So -2+6 is?

OpenStudy (ibbutibbu.):

oh i thought it was a negative 2

OpenStudy (ibbutibbu.):

ok so yeah the answer is 4

myininaya (myininaya):

So the ordered pair from when he is 2 meters from the goal is (2,-2+6)

OpenStudy (ibbutibbu.):

(2,4)?

myininaya (myininaya):

your question is to find the ordered pair when he is p meters away from the goal

myininaya (myininaya):

so instead of plugging in 2 we are going to plug in p

OpenStudy (ibbutibbu.):

so (2, p+6)

myininaya (myininaya):

well no remember ordered pairs look like (x,f(x)) and x is no longer 2 because we are looking at when he is p meters away from the goal and also your negative is missing on your p in the y part

OpenStudy (ibbutibbu.):

O_O what

myininaya (myininaya):

Why did you put x is 2 in that ordered pair? And what happen to the negative?

myininaya (myininaya):

Remember f(x)=-x+6 not x+6

myininaya (myininaya):

you are replacing x with p only

OpenStudy (ibbutibbu.):

ohkay so.. umm.. i have no clue

OpenStudy (ibbutibbu.):

sorry..

myininaya (myininaya):

you said when he is 2 meters away we have (2,-2+6) if you can do that then you can find the ordered pair for when he is p meters away Just replace the x's in (x,f(x))

myininaya (myininaya):

That is all you did for x=2 I don't see how x=p is any different that what we did for x=2

OpenStudy (ibbutibbu.):

So.. (2,-p+6)??

myininaya (myininaya):

again we are not replacing p with 2

myininaya (myininaya):

you got the y-coordinate correct but for some reason you are replacing the coordinate with 2

myininaya (myininaya):

x-coordinate with 2*

myininaya (myininaya):

Look at 2 meters away we have (2,-2+6) At 3 meters away we have (3,-3+6) At 5 meters away we have (5,-5+6) At p meters away we have ( ? , -p+6)

myininaya (myininaya):

what should the ? be

OpenStudy (ibbutibbu.):

p?

myininaya (myininaya):

yes at p meters away we have -p+6 amount of hoops made so the ordered pair is (p,-p+6)

OpenStudy (ibbutibbu.):

Oh and that is the answer?

myininaya (myininaya):

What does the question ask?

OpenStudy (ibbutibbu.):

write an ordered pair

myininaya (myininaya):

Is (p,-p+6) an ordered pair?

OpenStudy (ibbutibbu.):

yes

myininaya (myininaya):

ok

OpenStudy (ibbutibbu.):

So it IS my answer right?

myininaya (myininaya):

Only if you are convinced it is? If you aren't convinced then what are you confused about?

OpenStudy (ibbutibbu.):

nevermind, can you pleez help me with one meore :)

myininaya (myininaya):

I have to check on someone else since he called me during this but I might be able to take a look in a few minutes we will see

OpenStudy (ibbutibbu.):

Ok i will attach it, so if you ever get time to pleez look at it. and ty:)

OpenStudy (ibbutibbu.):

Wow. SO many ppl came to this question at once

OpenStudy (loser66):

The best way to not be abused by numbers is abusing them back, hehehe

OpenStudy (rational):

lol whats with the name of the attachment i_hate_math

OpenStudy (ibbutibbu.):

I just need help with the actual question, i dont need critisizm

OpenStudy (rational):

btw one question per post is the best way to get quick help as new questions get bumped so that other people see and assist u

OpenStudy (ibbutibbu.):

ok i will open it in anodder question